a block of 1.5 kg is pushed up a rough 25 degrees inclined plane by a constant force of 15 newton acting parallel to the inclined plane. if the coefficient of the kinetic friction is 0.3 and the block is displaced by a distance 10cm along the inclined plane, calculate the work done against the force of gravity and against the force of friction


1
Expert's answer
2021-02-25T11:26:21-0500

a) Work done against the force of gravity can be found as follows:


"W_g=mgh."

We can find "h" from the geometry:


"sin\theta=\dfrac{h}{d},""h=dsin\theta."

Finally, we get:


"W_g=mgdsin\theta=1.5\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot0.1\ m\cdot sin25^{\circ}=0.62\ J."

b) We can find the work done against the force of friction as follows:


"W_{fr}=F_{fr}d=mgd \mu cos\theta,""W_{fr}=1.5\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot0.1\ m\cdot0.3\cdot cos180^{\circ}=-0.441\ J."